if you mix 10 liters of 50% acid solution with 30 liters of 20% acid solution, then what is the concentration of acid in teh resulting mixture?
Answer by nerdybill(7384)

if you mix 10 liters of 50% acid solution with 30 liters of 20% acid solution, then what is the concentration of acid in teh resulting mixture?
.
Let x = the concentration of the resulting mixture
then
.50(10) + .20(30) = x(10+30)
5 + 6 = x(40)
11 = x(40)
11/40 = x
0.275 = x
or
27.5% = x